OpenStudy (anonymous):

limit x-> infinity (e^3x-5x)^1/x) please help me on this... i got infinity over infinity so i used l'hopitals rule... but i couldn't complete the solution... my friend said its e^3, but i dont know how to get it...

geerky42 (geerky42):

Just to be clear, is it \(\Large\displaystyle \lim_{x \rightarrow\infty} (e^{3x}-5x)^{\frac{1}{x}}\)?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

my solution is up to.. ln y = lim x-> infinity (3e^3x - 5)/(e^3x - 5x) .. what will i do next?

zepdrix (zepdrix):

So in it's current state, it's approaching the indeterminate form: \(\Large\rm \infty^0\) So we uhh, rewrite our expression using log and exponential,\[\LARGE e^{\ln\lim_{x \rightarrow\infty} (e^{3x}-5x)^{\frac{1}{x}}}\]Ignore the e for now, applying rules of logs,\[\Large\rm \lim_{x\to\infty}\frac{1}{x}\ln(e^{3x}-5x)\]Then,\[\Large\rm \lim_{x\to\infty}\frac{\ln(e^{3x}-5x)}{x}\]Now it's approaching \(\Large\rm \dfrac{\infty}{\infty}\)So this is where you applied L'Hopital I assume?

zepdrix (zepdrix):

\[\Large\rm \lim_{x\to\infty}\frac{3e^{3x}-5}{e^{3x}-5x}\]Woops.

OpenStudy (anonymous):

@zepdrix yeah...

zepdrix (zepdrix):

So we're STILL getting the indeterminate form \(\Large\rm \dfrac{\infty}{\infty}\), yes? So L'Hop again! :)

OpenStudy (anonymous):

i got it! thanks for the info :)

zepdrix (zepdrix):

cool c: just had to L'Hop a ton of times for this one lol

OpenStudy (anonymous):

what will be the limit, if i use the same function, as x->0...?

zepdrix (zepdrix):

Hmm well let's see.. everything would be the same up to this point:\[\Large\rm \lim_{x\to\infty}\frac{3e^{3x}-5}{e^{3x}-5x}\]Before this point you would be getting the indeterminate form 0/0 allowing you to apply L'Hop as you had before. As x->0, this expression is approaching:\[\Large\rm \lim_{x\to\infty}\frac{3e^{3x}-5}{e^{3x}-5x}=\frac{3(1)-5}{(1)-0}\]The exponentials are approaching 1 as x approaches 0, yah? So looks like we get uhhhhhhh, e^{-2} maybe? I dunno I'll try to confirm it on Wolfram or something.

zepdrix (zepdrix):

Wolfram agrees! Yay!

zepdrix (zepdrix):

Ah I didn't change the subscript to x->0 on the limits! My bad.

OpenStudy (anonymous):

thanks :)